A car starts from the rest covers a distanc 500m while accelerating uniformly at 10m/s2.


1
Expert's answer
2022-02-23T08:30:15-0500

Given:

"v_0=0"

"d=500\:\rm m"

"a=10\:\rm m/s^2"


The final velocity

"v=\sqrt{v_0^2+2ad}=\sqrt{0^2+2*10*500}=100\:\rm m/s"